OpenStudy (anonymous):

please help me solve the initial value problem y' = x^3(1-y) , y(0) = 3

OpenStudy (anonymous):

just separate the variables \[\Large \frac{dy}{1-y}=x^3dx\] integrate both sides can you do this now?

OpenStudy (anonymous):

one sec, lemme give it a go

OpenStudy (lgbasallote):

would be fun to use laplaces here :D

OpenStudy (anonymous):

nah, have to brush up on my separable equations

OpenStudy (anonymous):

it would be great with Laplace:P

OpenStudy (anonymous):

@IStutts can you do this or should i guide you ?

OpenStudy (lgbasallote):

it seems it can also be solve by linear DE

OpenStudy (anonymous):

yes you can solve this with integrating factor as well.

OpenStudy (anonymous):

A little guidance please. should I got Ln|1-y|?

OpenStudy (lgbasallote):

that was soooo close...

OpenStudy (anonymous):

yes you should get -ln(1-y) :P

OpenStudy (anonymous):

BLAST!

OpenStudy (lgbasallote):

haha

OpenStudy (anonymous):

then on the other side i would get x^4/4?

OpenStudy (anonymous):

+c

OpenStudy (anonymous):

you forgot the constant C

OpenStudy (anonymous):

yes !!

OpenStudy (anonymous):

haHAAAAA

OpenStudy (anonymous):

pellete.... so then i e to both of those equations? e^(-ln|1-y|) = e^(x^4/4)*e^c

OpenStudy (anonymous):

which cancels out the ln on the left side?

OpenStudy (anonymous):

it will but it will be okay if you write the left side as ln(1-y)^-1 e^ln(1-y)^-1 1/(1-y) !!

OpenStudy (anonymous):

not quite sure why you did that

OpenStudy (anonymous):

so in the end, i get y = 2e^(-x^4/4) +1?

OpenStudy (anonymous):

sorry i was busy with other question

OpenStudy (anonymous):

ok, i am posting a new one that i need assistance through

OpenStudy (anonymous):

ok post it.

OpenStudy (anonymous):

BOOM!